# The T2K Collaboration Signal selection and model-independent extraction of the neutrino neutral-current single cross section with the T2K experiment K. Abe

### v2

@@ V.1 Model Variations
Biases of at most 5% the uncertainty are observed when performing normalization shifts of several backgrounds, including variations of OOFV, CC, CC, NCX, CCQE, CC0, NCE, NC-other and NC1 with protons above 200 MeV/, corresponding to IDs 0-8. Particularly relevant is the result of increasing the weight of NC1 background events with protons, which supports the signal definition choice earlier described in Sec. III.3. The applied Np increase corresponds to the prediction difference between NEUT and GENIE for NC events with protons regardless of their momentum. It is worth noting that an analogous test with an alternative signal definition including exclusively protonless NC events yielded a bias of approximately 30% of the measurement uncertainty, demonstrating the importance of an adequate signal definition.
Scaling tests for events out of the RoI (0.5 and GeV/), corresponding to high angle events and low momentum events translate into biases of about 10% of the uncertainty, and are associated to IDs 9 and 10.
Variations of background events with low (VA) and high (VA) vertex activity, corresponding to IDs 11 and 12, yield biases of about 5% of the uncertainty. The VA variable denotes the summed light yield in a 55 volume of FGD1 bars around the reconstructed vertex, and is highly sensitive to the presence of low-momentum hadrons.
A re-scaling of the CCQE events based on their true value, was done in three ways: 1) an overall normalization of all events with low , i.e. all those events below 0.25 GeV2/c2; 2) a parametric modification of those events using 5 normalization parameters for GeV2/c2 taken in steps of GeV2/c2, and with values corresponding to the pre-fit and 3) post-fit values for those model parameters used in the oscillation analysis of T2K in Ref Abe et al. (2023a). In all cases observed biases are about 5% of the measurement uncertainty. These tests have associated IDs 13-15.
Positive and negative normalization shifts of the signal cross section, corresponding to IDs 16 and 17, lead to negligible biases.
@@
The observed biases in all cases are a small fraction of the measurement uncertainty supporting the flexibility of the fitted model and the adequacy of the size of the prior uncertainties.
Table 3: Summary of fake data studies results. All expectations use Monte Carlo simulations normalized to the data POT. The ”Ratio” column shows the ratio of the integrated cross section from each study to that obtained from an Asimov fit using nominal NEUT event rates.
@@ VI Cross Section Result
The data fit has a , divided in and . The associated p-value is 0.11, supporting a statistically reasonable goodness-of-fit.
Figure 11: Differential cross section and post-fit uncertainty expected from the Asimov fit of the NEUT simulation compared to the measured data result.Figure 12: Differential cross section result and post-fit uncertainty compared to the Asimov fit of the NEUT simulation. The x-axis error bar covers one standard deviation of the T2K neutrino flux around its mean energy of 0.70 GeV.

### v3

@@ V.1 Model Variations
Biases of at most 5% the uncertainty are observed when performing normalization shifts the most important backgrounds for the analyses, including variations of OOFV, CC, CC, NCX, CCQE, CC0, NCE, NC-other and NC1 with protons above 200 MeV/, corresponding to IDs 0-8. Particularly relevant is the result of increasing the weight of NC1 background events with protons, which supports the signal definition choice earlier described in Sec. III.3. The applied Np increase ( corresponds to the prediction difference between NEUT and GENIE for NC events with protons regardless of their momentum. It is worth noting that an analogous test with an alternative signal definition including exclusively protonless NC events yielded a bias of approximately 30% of the measurement uncertainty, demonstrating the importance of an adequate signal definition.
Scaling tests to check the impact of a varied normalization for the events out of the RoI (0.5 and GeV/), were we do not report the measurement due to low selection purity and efficiency, translate into biases of about 10% of the magnitude of the assessed uncertainty, and are associated to IDs 9 and 10.
Variations of background events with low (VA) and high (VA) vertex activity, corresponding to IDs 11 and 12, yield biases of about 5% of the magnitude of the uncertainty. The VA variable denotes the summed light yield in a 55 volume of FGD1 bars around the reconstructed vertex, and therefore this test checks for the impact of mis-modeling in the predictions of low-momentum hadrons.
A re-scaling of the CCQE events based on their true value was done in three ways to check the impact of -dependent mis-modeling hinted by several experiments: 1) an overall normalization of all events with low , i.e. all those events below 0.25 GeV2/c2; 2) a parametric modification of those events using 5 normalization parameters for GeV2/c2 taken in steps of GeV2/c2, and with values corresponding to the pre-fit and 3) post-fit values for those model parameters used in the oscillation analysis of T2K in Ref Abe et al. (2023a). In all cases observed biases are about 5% of the measurement uncertainty. These tests have associated IDs 13-15.
Positive and negative normalization shifts of the signal cross section, corresponding to IDs 16 and 17, lead to negligible biases, confirming the adequate fitter behavior to capture signal variations.
@@
While Table 3 summarizes the results in terms of the integrated cross-section values, we also looked for indications of bias in differential bins for each study. The observed biases in all cases are a small fraction of the measurement uncertainty supporting the flexibility of the fitted model and the adequacy of the size of the prior uncertainties.
Table 3: Summary of fake data studies results. All expectations use Monte Carlo simulations normalized to the data POT. The ”Ratio” column shows the ratio of the integrated cross section from each study to that obtained from an Asimov fit using nominal NEUT event rates. The post-fit corresponds to Eq. 2 at the best-fit-point.
@@ VI Cross Section Result
Data unbinding was performed in a two-stage process. Firstly, the pre-fit event rate in the sidebands was checked to be in qualitative agreement with data and a fit to the side-bands using nominal predictions for the signal sample was performed. No nuisance parameters (detector, flux or cross-section) were observed to be in significant tension with their nominal values. Secondly, the signal sample was also unblinded, leading to the final data fit and result that we report next.
The data fit has a , divided in and . The associated p-value is 0.11, supporting a statistically reasonable goodness-of-fit. The largest contribution to is the statistical term, which is in a similar level of agreement in the signal sample and the sidebands. In particular, the statistical in the signal, EPID, AddTrk and TPID samples corresponds to: , , , and .
Figure 11: Differential cross section from NEUT v5.4.0 compared to the measured differential data result.Figure 12: Integrated cross section from NEUT v5.4.0 compared to the measured integrated data result. The x-axis error bar covers one standard deviation of the T2K neutrino flux around its mean energy of 0.70 GeV.
@@
Additional robustness tests on the final result are provided in Appendix A.
@@ VII Conclusions
The measurement results have an associated data release that can be found in Ref. T2K Collaboration (2025).
@@ Acknowledgments
We show in Fig. 13 the observed variations in the final cross-section results when using alternative cross-section model parameters during the unfolding. The alternative tests are as follows:
1.
Disabled: We switch off completely the cross-section model parameters. Namely, we only use detector and flux parameters during unfolding.
2.
Norm-Only: All shape parameters (, 2p2h shape, and , and FSI parameters) are switched off.
3.
Low-Q2: All the low- normalization parameters are switched off.
4.
No OOFV: All the OOFV normalization parameters are switched off.
Even in the most extreme case, where the cross-section model is fully disabled, the alternative result is well within one sigma of the reported cross-section result for each differential bin. In all cases the pull between the NEUT 5.4.0 prediction and the final result is much larger than the observed variations for the alternative cross-section models, supporting the robustness of the result. Furthermore, fits where certain sub-sets of cross-section models are disabled (Norm-Only, Low-, No OOFV), which directly affect the most abundant categories of selected events in the signal sample and the sidebands, lead to minimal changes in the reported result.
Figure 13: Differentiable cross section for NEUT 5.4.0 and the data result, as well as the data results for alternative cross-section model parametrization choices.
Appendix B Binning
Table 4: Binning in true kinematic space. Bins 0, 1, 11 and 12 cover the kinematics space out of the region of interest.
Index
min
max
min p (MeV/)
max p (MeV/)
0
-1
0.5
0
30000
1
0.5
1
0
200
2
0.5
0.7
200
600
3
0.7
0.8
200
600
4
0.5
0.8
600
1000
5
0.8
0.9
200
400
6
0.8
0.9
400
600
7
0.8
0.9
600
1000
8
0.9
1.0
200
400
9
0.9
1.0
400
600
10
0.9
1.0
600
1000
11
0.5
0.9
1000
30000
12
0.9
1.0
1000
30000
Table 5: Binning in reconstructed kinematic space for the signal sample.
Index
min
max
min p (MeV/)
max p (MeV/)
0
-1.0
0.5
0
30000
1
0.5
0.8
0
200
2
0.8
1.0
0
200
3
0.5
0.6
200
600
4
0.6
0.7
200
350
5
0.6
0.7
350
600
6
0.5
0.8
600
1000
7
0.5
0.8
1000
30000
8
0.7
0.8
200
300
9
0.7
0.8
300
400
10
0.7
0.8
400
600
11
0.8
0.9
200
300
12
0.8
0.9
300
400
13
0.8
0.9
400
500
14
0.8
0.9
500
600
15
0.8
0.9
600
800
16
0.8
0.9
800
1000
17
0.8
0.9
1000
1500
18
0.8
0.9
1500
30000
19
0.9
1
200
300
20
0.9
1
300
400
21
0.9
1
400
500
22
0.9
1
500
600
23
0.9
1
600
700
24
0.9
1
700
850
25
0.9
1
850
1000
26
0.9
1
1000
1200
27
0.9
1
1200
1500
28
0.9
1
1500
1700
29
0.9
1
1700
2000
30
0.9
1
2000
2400
31
0.9
1
2400
3200
32
0.9
1
3200
30000
Table 6: Binning in reconstructed kinematic space for the EPID sample.
Index
min
max
min p (MeV/)
max p (MeV/)
33
-1.0
0.5
0
30000
34
0.5
1.0
0
300
35
0.5
0.7
300
600
36
0.5
0.7
600
30000
37
0.7
0.8
300
600
38
0.7
0.8
600
800
39
0.7
0.8
800
30000
40
0.8
0.9
300
500
41
0.8
0.9
500
600
42
0.8
0.9
600
800
43
0.8
0.9
800
1000
44
0.8
0.9
1000
1300
45
0.8
0.9
1300
30000
46
0.9
1
300
500
47
0.9
1
500
600
48
0.9
1
600
700
49
0.9
1
700
800
50
0.9
1
800
1100
51
0.9
1
1100
1500
52
0.9
1
1500
2000
53
0.9
1
2000
2500
54
0.9
1
2500
3500
55
0.9
1
3500
30000
Table 7: Binning in reconstructed kinematic space for the AddTrk sample.
Index
min
max
min p (MeV/)
max p (MeV/)
56
-1.0
0.6
200
30000
57
0.6
0.8
200
30000
58
-1
1.0
0
200
59
0.8
1.0
200
600
60
0.8
1.0
600
1500
61
0.8
1.0
1500
30000
Table 8: Binning in reconstructed kinematic space for the TPID sample.